OpenStudy (anonymous):

Given f(x) = 2x-3 , what is f(-5) ? A. –16 B. –13 C. 7 D. 30 @dan815 @tanya123

OpenStudy (texaschic101):

f(x) = 2x - 3 f(-5) = 2(-5) - 3 f(-5) = -10 - 3 f(-5) = -13

OpenStudy (texaschic101):

all you do is sub in -5 for every place there is an x